INAF queue observing, 2016 December 4 - 5

Observer: R. Carini, A. Marchetti
Telescope Operator: G. Bechetti
Support Astronomer: O. Kuhn on call
Instrument:

Summary

LUCI2 is not available

At sunset the sky is clear. There is wind coming from the north-east. We need to choose targets towards the south.

Night duration (12\xB0 el twilight) 18:10-6:00 ( 01:10-13.00 UT) = 11h 50m

Observing time: 9h 30m (80%)

Weather time loss: 0h 0m (0%)

technical time loss 2h 20m (20%)

Data Summary

Project_RUN | Objects | Filter | ExpTime | Notes |
Marziani B SDSS J0008-1039 H 18m seeing very variable. taken the last two images to complete the target, plus one due to bad seeing.
telluric Marzini B HIP1378 H    
Antoniucci ZXtau zj 0.2 h miss HK filter and ima
telluric Iovino HIP11809 HKspec    
Iovino M004XXLrev HKspec 1 h last frame with 1.4 < airmass <1.43 and 0.85'' < seeing< 1.00''
telluric Antoniucci HIP30393 zj    
Castellano ob. 17 MACJ0717 Uspec, V COMPLETED
Castellano ob. 18 MACJ0717 Uspec, V COMPLETED
Castellano ob. 19 MACJ0717 Uspec, V COMPLETED, but seeing oscillating. CHECK
photometric star SA101 blue:Uspec, g, red:V,r    
Bellazzini I 1437+12.9 blue:g,red:r 0.2h COMPLETED
Bellazzini J 147.0+07.1 blue:g,red:r 0.2h COMPLETED
photometric star SA104 blue:Uspec, g, red:V,r
DAlessandro Pal14 1 image in blue, 3 in red, during twiligth
CALIBRATION sky flats for LBC blue U_spec, g, red V, r
CALIBRATION bias for LBC

Details

MARZIANI B

01:10 We send the TELLURIC script for Marziani B target
  • seeing about 1"
  • acquisition luci1.20161205.0001-0008 BAD spectra (there was a problem in the telluric acquisition script like yesterday: wrong combination of filters, with OH and H no light passes!)
  • luci1.20161205.0009-0010 good telluric spectra
  • 01:30 seeing oscillating between 1.2" and 1,5"

01:38 we send the SCIENCE script for Marziani B target (only last two frames are needed to complete the target observation)
  • acquisition luci1.20161205.0011-0014
  • 01:54 seeing about 1.5"
  • 01:56 seeing 1.3"
  • 01:58 we cannot see the target inside the slit in the confirmatory image, and subtracting one frame from another to see it is incredibly slow. It looks like it crashed, we stop it and kill the window. We lose the observer panel. We open it again. The alignment information was not lost.
  • 02:01 we try again to subtract the sky to the thru-slit image. Now we see the target in the slit.

02:05 we take the target spectra luci1.20161205.0015-0017
  • 02:10the seeing is now 1.5", but it keeps oscillating between 1.2" and 1.8" in the last 10 minutes
  • 02:22 the seeing is still oscillating between 1.2" and 1.8"
  • we take an extra image because the seeing is very variable.

02:38 the seeing is very variable. we decide to run a backup target of Antoniucci
  • begin the acquisition of the telluric for V512Cep, but the star is direct into the wind, even if the target is good.
  • 02:41 reset the shell

02:45 acquire the target XZtAu
  • seeing 1.7 airmass 1.8
  • 02:50 the TO hast to fix the collimation
  • 02:59 send again the preset for XZtau luci1.20161205.0018-0021
  • 03:01 seeing 1''

03:10 begin science luci.0022-0025 zj filter is OK
  • the PI made a mistake in the script, images in HK filter are not good (luci.0026-0029).

03:23 the seeing is good and stable, we slew to Iovino target
  • acquisition of telluric luci.0030-0031
  • seeing 0.85''

03:36 engineering problem about the alingment
  • to fix the problem we have token some images with the AS with and without mask luci.0032-0075

04:13 problem resolved, we send the preset again for the telluric.
  • 04:14 the pointing is off, the TO fixes it.

04:17 we send again the telluric script 2015_2016_5_HK_tell_LUCI1.xml
  • 04:28 seeing 0.9''-1.00''
  • acquisition frames luci.0076-0078
  • science frames luci.0079-00786

04:47 send the science script 2015_2016_5_HK_LUCI1.xml
  • 04:51 problem: the mask can not move from the focus plane to the turnout position
  • 05:10: fix the problem. the name of the mask in the script was wrong

05:11 send again the script 2015_2016_5_HK_LUCI1.xml
  • seeing 0.7'''
  • frames of acquisition luci.0087-0090
  • frames os science luci.0091-009
  • 06:01 seeing 0.8''
  • 06:18 seeing 0.8''-0.9''
  • 06:32 last image of the script taken with 1.4 < airmass < 1.43 but seeing 0.85''-1.00''

06:38 run the telluric for Antoniucci XZTau
  • acquisition luci1.0103-0105
  • science luci1.0106-0107 zj filter
  • seeing 0.9'&

06:58 switch to LBC
  • 07:12 problem with the switch. the telescope have problem with the balance.LBC in the beam let balance the telescope

08:07 problem fixed
  • begin collimations
  • 08:09* error : do not authorize the cameras
  • 08:17* fixed the problem

08:07 problem fixed
  • begin collimations
  • 08:09* error : do not authorize the cameras
  • 08:17* fixed the problem

08:20 begin again the collimation: dohybrid
  • 08:26 copointing
  • 08:31 run science: Castellano_MACSJ0717_UV_17.ob
  • seeing 1.0'';
  • 08:58 seeing 0.7'';

09:14 copointing

09:19 run science Castellano_MACSJ0717_UV_18.ob. Seeing 0.8'';
  • 09:45 seeing 0.8'';
  • 09:56 seeing 1.0'';

10:01 dofpia

10:07 run science Castellano_MACSJ0717_UV_19.ob. Seeing begins to be very variable, with picks up to 1.2"
  • 10:26 seeing between 0.8'';-0.9'';
  • 10:39 the wind blows NW, the same direction of our target. We have to change target
  • during this ob the seeing has been very variable, the images should be checked one by one. Probably the first 8 images are not good
  • 10:46 seeing 1.1'';

10:49 we switch to photometric star SA101. Technical problem: lost communication between host and tracker in the blue channel. We try to solve the problem sending the preset again
  • 11:03 LBC got stuck. After calling the SA, we have to turn it off.
  • 11:14 we turn off LBC
  • 11:17 we turn on LBC again
  • 11:22 we run SA101_Sg_Vr

11:35 seeing is good, < 1''; we slew to Bellazzini I
  • 11:47 dofpia
  • 11:55 copointing

11:59 run science of Bellazzini I. Seeing is about 0.7';

12:13 we slew to Bellazzini target J
  • 12:14 dofpia. Seeing 0.7'';-0.8'';
  • 12:18* copointing

12:23 send science for Bellazzini target J. Seeing is 0.7''; *12:30 seeing 0.8'';-0.9'';

12:39 we slew to SA104 photometric star

12:49 we slew to D'Alessandro Pal4

12:51 dofpia

12:57 copointing

13:01 we send science for Dalessandro. *13:09 we are in twilight, we took 1 image for the blue and 3 for the red and then stopped the script

13:11 we start taking sky flats, first gr, then SV. There are clouds coming in. The flats are taken with clouds

-- %BUBBLESIG{AlidaMarchetti - 2016-12-05}%
 
Topic revision: r12 - 08 Dec 2016, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback